Instrument Log

GPS ARRAY              C1                 Central Facility, Lamont, OK
 

Date

Technician

Comments

7/14/10

CAM/JEM

GPS instruments are being decommissioned and need to be pulled from the field.  Lowered the 60M tower SE booms at 1445 GMT.  Removed the repeater Yagi and whip antennas, control box with two transceivers and power supplies, and all associated hardware and cables.  Transceivers MD# DGR-115H, SNs 910-2322 and 910-0824.  Raised booms and tower RTS at 1556 GMT.

7/14/10

CAM/JEM

GPS instruments are being decommissioned and need to be pulled from the field.  Powered the computer off in the optical trailer at 1400 GMT and removed all computer equipment, receivers and antennas from the trailer.  This included the transceiver for site 0N0W (SN 900-8926) and the transceiver for the repeater (MD# DGR-115H, SN 571-1108).

7/13/10

CAM/JEM

GPS instruments are being decommissioned and need to be pulled from the field.  Removed the following sites today:

1N1E pulled at 1820 GMT.  Only had the solar panel bracket.  Rest was stolen.

1N0W pulled at 1830 GMT.  Transceiver SN 900-8891

1N1W pulled at 1840 GMT.  No transceiver or GPS antenna was installed here.

1N2W pulled at 1910 GMT.  Transceiver SN 900-8893

2N3W pulled at 1920 GMT.  Transceiver SN 900-8922

2N0W pulled at 2000 GMT.  Transceiver SN 900-8930

2N2W pulled at 2010 GMT.  Transceiver SN 900-8899

2N1W pulled at 2020 GMT.  Transceiver SN 902-7976

2N1E pulled at 2040 GMT.   Transceiver SN 900-8923

7/12/10

CAM/JEM

GPS instruments are being decommissioned and need to be pulled from the field.  Removed the following sites today:

2S0W pulled at 1530 GMT.  Transceiver SN 902-7981

2S1W pulled at 1610 GMT.  Transceiver SN 902-9049

2S2W pulled at 1630 GMT.  Transceiver SN

2S3W pulled at 1645 GMT.  Transceiver SN 902-7984

2S1E pulled at 1715 GMT.   Transceiver SN 902-7994

1S0W pulled at 1910 GMT.  Transceiver SN 901-3418

1S1W pulled at 1925 GMT.  Transceiver SN 900-8950

1S2W pulled at 1940 GMT.  Transceiver SN 902-7995

1S3W pulled at 1955 GMT.  Transceiver SN 900-7034

0N2W pulled at 2010 GMT.  Transceiver SN 901-3418

0N1W pulled at 2045 GMT.  Transceiver SN 900-8689

0N1E pulled at 2100 GMT.   Transceiver SN 900-0442

10/28/09

KT/J. Braun

At 1516 GMT performed maintenance on the GPS units with instrument mentor. All sites were visited and some repairs were made.

 

2S1E – The antenna was removed and replaced. System was left tracking.

2S1W – The antenna was removed and replaced. System would still not track and possibly needs the receiver to be replaced.

1S3W – Replaced the power cable. Reconfigured the radio modem.

1S1W – Removed and replaced the antenna. System was left tracking.

0N1E – Removed and replaced the antenna. System was left tracking.

1N2W – Installed Transmitter and antenna cable.

1N0W – Instrument needs new antenna. Reconfigured the radio modem.

 

Instrument mentor would like to upgrade the antenna poles from fiber glass to

metal. The tops of the GPS antenna’s need to be resealed with all weather caulking. Instrument mentor will be sending out spare antenna’s that need to be changed out with other sites.                                    

10/26/09

CAM

Checked status of all sites.  All sites appear to be operating normally according to the status of the LEDs with the following exceptions.

2N3W: CD solid red, CTS slowly flashing red, TX shows no activity.

1N0W: CD solid red, CTS slowly flashing red, TX shows no activity.

1N2W: Receiver has been returned to John Braun.

1N1W: Receiver has been returned to John Braun.

1S3W: Receiver has been returned to John Braun.

9/23/09

CAM

John Braun requested that the lightning suppressor be removed from the GPS antenna input at site 0N1E.  Removed lightning suppressor from both the GPS and YAGI antennas at 2210 GMT.  This is part of T/S efforts to establish GPS link.

9/22/09

CAM

Performed the following maintenance as requested by John Braun.

0N1E: Replaced GPS antenna at 1922 GMT

0N2W: Replaced GPS antenna at 1939 GMT. 

2N1E: Replaced GPS antenna at 2007 GMT. 

2S1E: Replaced GPS antenna cable at 1905 GMT.

2N0W: Replaced GPS antenna cable at 1958 GMT. 

Cleaned light corrosion of the cable/antenna connectors at each site.  LED status indicated good operation at all the above sites before and after maintenance.

4/9/09

CAM

Checked status of the 8 sites north of the CF on 1N and 2N roads.  Status as follows:

1N0W: All LEDs were dimly lit green.  Power cycled receiver at 1930 GMT.  All LEDs came back on lit bright, solid red.  Status is not operational.

1N1W:  Receiver was removed previously and returned to John Braun.

1N2W:  Receiver was removed previously and returned to John Braun.

2N1E:  Status O.K.

2N0W: Status O.K.

2N1W: CD lit solid red.  CTS flashing red.  No TX.  Found Yagi antenna cable cut in two.  Replaced cable with one from site 1N2W (no receiver) at 2038 GMT.  LEDs now indicate site is operational.  Will repair cable and return to 1N2W.

2N2W:  Status O.K.

2N3W:  CD lit solid red.  CTS flashing red.  No TX.  Power cycled receiver at 1955 GMT with no change in status.

4/8/09

CAM

Checked status of the 12 sites on 0N, 1S and 2S roads.  All sites were operational with the exception of 0N1E and 1S3W.  Site 0N1E LEDs were flashing alternately between all red and all green.  Power cycled receiver at 1930 GMT.  Status of LEDs returned to normal.  Site 1S3W receiver has been returned to John Braun and is not in service.  Will check status of remaining 8 sites tomorrow.

1/22/09

CAM

Site 0N2W.  GPS unit had no power.  Battery wiring had been destroyed by rodents.  Removed old wiring and replaced with new.  Restored power and RTS at 2130 GMT.

1/21/09

CAM

Checked operational status of all sites today.  Status of all sites unchanged from previous visit, with the exception of 2N3W and 0N2W.  Site 2N3W now appears to be operating as it should. Previously CD light was solid red.  Site 0N2W has no power due to rodents destroying battery wiring.  Will repair tomorrow.

1/21/09

CAM

Site 1S2W.  The insulation on the battery cables had been stripped by rodents.  The exposed cables were in danger of shorting together.  Disconnected battery cables at 2015 GMT.   Wrapped bare areas with electrical tape.  Reconnected cables and RTS at 2030 GMT.

8/18/08

CAM

John Braun requested the power cables be replaced at stations 2N3W and 1N0W.  Replaced power cables at 2N3W and 1N0W at 1823 and 1845 GMT, respectively.  New cables do not appear to have made a difference.  Receivers are getting power, but CD is still solid red and CTS is flashing red slowly at both sites.

8/18/08

CAM

John Braun requested the GPS receivers be pulled from stations 1S3W and 1N2W and returned to him for reconfiguration.  Pulled GPS receivers from stations 1S3W and 1N2W at 1758 and 1808 GMT, respectively.  Receivers will be shipped to John today via FedEx.

8/18/08

CAM

John Braun requested the GPS antennas be replaced at stations 2S1E, 1S0W and 2N0W.  Replaced 2S1E GPS antenna at 1732 GMT, 1S0W GPS antenna at 1745 GMT, and 2N0W GPS antenna at 1833 GMT.

8/12/08

CAM

Received 10 of the 11 receivers back from John Braun and need to install.

Installed 2S1E receiver (SN 902-7994) at 1410 GMT.  No lights came on.  Battery voltage was down to 9V.  Checked solar panel output and found it was only 9V.  Replaced solar panel and regulator with spare.  Also replaced battery.  RTS at 2000 GMT.  Initially CD was alternating between green and red, but was green at last check.

Installed 1S0W (SN 901-3418) at 1422 GMT.  Status O.K.

Installed 1S1W (SN 900-8950) at 1432 GMT.  Status O.K.

Installed 1S2W (SN 902-7995) at 1442 GMT.  Status O.K.

Installed 0N1W (SN 900-8689) at 1505 GMT.  Status O.K.

Installed 1N0W (SN 900-8898) at 1518 GMT.  CD solid red.  No TX activity.

Installed 2N1E (SN 900-8923) at 1540 GMT.  Status O.K.

Installed 2N0W (SN 900-8930) at 1546 GMT.  Status O.K.

Installed 2N3W (SN 900-8922) at 1555 GMT.  CD solid red.  No TX activity.

Installed 0N1E (SN 900-0442) at 1630 GMT.  Status O.K.

7/14/08

KT

Removed the 0N1E GPS receiver at 1450 GMT. SN# 902-7978

Removed the 0N1W GPS receiver at 1510 GMT. SN# 900-8689

Removed the 1N0W GPS receiver at 1530 GMT. SN# 900-8898

Removed the 2N1E GPS receiver at 1543 GMT. SN# 900-8923

Removed the 2N0W GPS receiver at 1555 GMT. SN# 900-8930

Removed the 2N3W GPS receiver at 1610 GMT. SN# 900-8922

Removed the 1S0W GPS receiver at 1630 GMT. SN# 901-3418

Removed the 1S1W GPS receiver at 1645 GMT. SN# 900-8950

Removed the 1S2W GPS receiver at 1700 GMT. SN# 902-7995

Removed the 2S1E GPS receiver at 1743 GMT. SN# 902-7994

Removed the 1N1W GPS receiver at 1820 GMT. SN# 900-8903

All receiver model numbers DGR-115WGI.  Shipped receivers to John Braun to be reconfigured.

6/26/08

CAM

Performed maintenance on system 2N1W:

Status before maintenance: Not operational.  No lights lit.

Rodents had chewed insulation off battery cables and cables were shorted out.  Cleaned rodent nest from battery box.  Replaced battery and repaired cables.  RTS at 2105 GMT.  Status after maintenance:  Lights indicate good operation.

6/26/08

CAM

Performed maintenance on system 2S1E:

Status before maintenance: Not operational.  No lights lit.

Battery voltage measured 4.0V.  Cleaned rodent nest from battery box.  Replaced battery and repaired cables.  RTS at 2020 GMT. 

Status after maintenance:  Lights indicate bad operation.  CD is solid red.  CTS is flashing red.  TX status is uncertain.  Cannot tell if it is flashing or not.

6/23/08

CAM

Performed maintenance on system 2S1W:

Status before maintenance: Lights indicated system was operational. 

However, the antenna mast was broken and laying on the ground. Cleaned rodent nest from battery box.  Replaced battery.  Replaced broken antenna mast with spare.  RTS at 2100 GMT. 

Status after maintenance:  Lights indicate good operation. 

6/23/08

CAM

Performed maintenance on system 2N2W:

Status before maintenance: Not operational.  No lights lit.

Mice had chewed through the battery cables.  Cleaned rodent nest from battery box.  Replaced battery and repaired cables.  RTS at 1940 GMT. 

Status after maintenance:  Lights indicate good operation. 

6/23/08

CAM

Performed maintenance on system 2N3W:

Status before maintenance: Not operational.  No lights lit.

Mice had chewed through the battery cables.  The antenna mast was broken and laying on the ground. Cleaned rodent nest from battery box.  Replaced battery and repaired cables.  Replaced broken antenna mast with spare.  RTS at 1900 GMT.  Status after maintenance:  Lights indicate bad operation.  CD is solid red.  CTS is flashing red.  TX has no activity.

6/2/08

CAM

Visited system 2N3W and found the antenna mast was laying on the ground.  The threaded end of the mast is broken.  Mast needs to be replaced.

6/2/08

CAM

Performed maintenance on system 2N0W:

Status before maintenance:  Lights indicate bad operation.  All are solid green.

Mice had removed insulation from battery cables.  Cables were shorted out. Cleaned rodent nest from battery box.  Replaced battery and repaired cables.   RTS at 2050 GMT. 

Status after maintenance:  Lights indicate bad operation.  All are solid green.

6/2/08

CAM

Performed maintenance on system 2N1E:

Status before maintenance:  Not operational – no lights.

Mice had chewed through both battery cables.  Cleaned rodent nest from battery box.  Replaced battery and repaired cables.   RTS at 2015 GMT. 

Status after maintenance:  Lights indicate bad operation.  All are solid green.

5/30/08

CAM

Performed maintenance on system 2S3W:

Status before maintenance:  Not operational – no lights.

Mice had chewed through the positive battery cable.  Cleaned rodent nest from battery box.  Replaced battery and repaired cables.  The GPS antenna connector is sheared off at the receiver.  Plugged the connector in and zip tied in place.  Will see about replacing with a spare.  RTS at 2042 GMT. 

Status after maintenance:  Lights indicate good operation.

5/30/08

CAM

Performed maintenance on system 2S2W:

Status before maintenance:  Lights indicated good operation.

Cleaned rodent nest from battery box.  Replaced battery. Taped sections of power wires where rodents had chewed insulation off.  RTS at 2010 GMT.

Status after maintenance:  Lights indicate good operation.

5/30/08

CAM

Visited system 2S1W and found the antenna mast was laying on the ground.  The threaded end of the mast is broken.  Mast needs to be replaced.

5/30/08

CAM

Performed maintenance on system 2S0W:

Status before maintenance:  Lights indicated good operation.

Cleaned rodent nest from battery box.  Replaced battery. Cut off bad sections of power wires where rodents had chewed insulation off.  Terminated cable ends.  RTS at 1940 GMT.

Status after maintenance:  Lights indicate good operation.

4/21/08

CAM

Performed maintenance on system 1N0W:

Status before maintenance:  Lights indicate bad operation.   CD and CTS solid red.  TX may be solid red too, but hard to tell for sure.

Replaced battery.   RTS at 2000 GMT.

Status after maintenance:  Lights indicate bad operation.   CD and CTS solid red.  TX may be solid red too, but hard to tell for sure.

4/21/08

CAM

Performed maintenance on system 1N1W:

Status before maintenance:  Lights indicated good operation.

Replaced battery.  RTS at 1935 GMT.

Status after maintenance:  Lights indicate good operation.

4/21/08

CAM

Performed maintenance on system 1N2W:

Status before maintenance:  Lights indicated good operation.

Replaced battery.  RTS at 1915 GMT.

Status after maintenance:  Lights indicate good operation.

4/21/08

CAM

Performed maintenance on system 1S3W:

Status before maintenance:  Not operational – no lights.

Mice had chewed through the battery cables.  Replaced battery and repaired cables.  Aligned antenna with 60M tower.   RTS at 1840 GMT.

Status after maintenance:  Lights indicate bad operation.   CD and CTS solid red.  TX may be solid red too, but hard to tell for sure.

4/21/08

CAM

Performed maintenance on system 1S2W. 

Status before maintenance:  Not operational – no lights.  

Mice had chewed through the battery cables.  Replaced battery and repaired cables.  RTS at 1810 GMT.

Status after maintenance:  Lights indicate good operation.

4/16/08

CAM

Need to install new ring terminals on the battery cables at 3 of the sites that previously had new batteries installed.  Installed new ring terminals at site 0N1W at 1955 GMT, 0N2W at 2015 GMT and 0N1E at 2035 GMT.

4/16/08

CAM

Performed maintenance on system 1S1W.  System appeared to be operational upon arrival to site as evident by lights flashing.  Battery is scheduled to be changed out.  Removed the old battery and cleaned mouse nest from battery box.  Cut the wires back to remove sections that were damaged by rodents.  Solar panel output measured ok at 19.0 VDC.  Old battery voltage was 13.2 VDC.  Replaced battery with new one.  System returned to service at 1930 GMT.

4/14/08

CAM

Performed maintenance on system 1S0W.  System appeared to be operational upon arrival to site as evident by lights flashing.  Battery is scheduled to be changed out.  Removed the old battery and cleaned mouse nest from battery box.  Solar panel output measured ok at 18.5 VDC.  Old battery voltage was 13.0 VDC.  Replaced battery with new one.  System returned to service at 1930 GMT.

4/14/08

CAM

Performed maintenance on system 0N2W.  System appeared to be operational upon arrival to site as evident by lights flashing.  Battery is scheduled to be changed out.  Removed the old battery and cleaned mouse nest from battery box.  Cut the wires back to remove sections that were damaged by rodents.  Solar panel output measured ok at 19.2 VDC.  Old battery voltage was 13.3 VDC.  Replaced battery with new one.  System returned to service at 1855 GMT.

4/14/08

CAM

Performed maintenance on system 0N1E.  System appeared to be operational upon arrival to site as evident by lights flashing.  Battery is scheduled to be changed out.  Removed the old battery and cleaned mouse nest from battery box.  Cut the wires back to remove sections that were damaged by rodents.  Solar panel output measured ok at 19.2 VDC.  Old battery voltage was 13.2 VDC.  Replaced battery with new one.  Aligned antenna with 60M tower and tightened base.  System returned to service at 1820 GMT.

4/10/08

CAM

Performed maintenance on system 0N1W.  System was not operational upon arrival to site as evident by no status light activity.  The solar panel has a hole through it from a rifle shot.  The bullet also nicked the solar panel controller box and cracked the lid.  Solar panel is still operational with 20VDC output.  Covered the holes in solar panel and controller box with tape. Found the battery box packed full with rats nest.  The battery housing had been chewed away significantly by rats around the positive post.  Battery cables were also chewed through.  Battery voltage was 11VDC.  Cut the wires back to remove bad sections.  Cleaned out the battery box and replaced the battery with a new one.  System returned to service at 1955 GMT.  Contacted John Braun, who confirmed system was now transmitting data.

3/10/08

PAD

Mentor request repeater on 60M tower and transceiver located in Optical trailer power cycled.  At 1427 GMT, power cycled repeater on side A of 60M tower.  At 1435 GMT, power cycled wireless data transceiver located in Optical trailer.  RTS at 1442 GMT.

7/26/07

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1333 GMT.  Raised booms and RTS at 1920 GMT.

6/14/05

MR

The spare compressor in the optical trailer has tripped a breaker causing the computer on the GPS array to be shut down.  Unplugged the compressor and reset the breaker.  Notified facilities about the compressor.  Restarted the GPS array computer at 1535 GMT.  RTS at 1545.

9/27/04

CAM

SE booms need to be lowered for 25m and 60m T/H calibrations.  Lowered booms at 1438 GMT.  Raised booms and RTS at 2056 GMT.

8/13/04

PAD

Instrument mentor notified that GPS network stopped collecting data on 8/11/04 at 0830 GMT.  Severe weather passed through area at that time.  Mentor requested equipment power cycled.  Power cycled instruments on Side A, (SE side) of tower at 60m level at 1603 GMT.  Instrument mentor notified GPS system returned to normal.  RTS at 1616 GMT.

4/15/04

CM/PD

SE booms need to be lowered for 25M MFR maintenance.  Lowered booms at 2104 GMT.  Raised booms and RTS at 2151 GMT.

10/24/03

PAD

At 1340 GMT, GPS Array battery backup was chirping.  Computer was down.  Found breaker number 4 in the optical trailer had been tripped.  Reset breaker at 1348 GMT. Computer restarted.  Notified instrument mentor.  RTS at 1355 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

1/30/03

JEM

The 60M Southeast temperature has been too high.  There is no problem with the GPS.  Corrective maintenance is being done on the 60M T/H.  Lowered Southeast 25M and 60M booms @1655 GMT.  RTS @1810 GMT.

10/31/02

CM/MR

SE booms lowered from 1445 to 1618 GMT for 25MFR and 60 T/H maintenance.

10/29/02

CAM

SE booms lowered from 2135 GMT to 2220 GMT for 25 and 60M T/H maintenance.

10/28/02

CAM

The SE 25 and 60M T/H sensors are reading –9999’s or 0’s.  Lowered the SE booms slightly at 1731 GMT and then raised to reseat connections.  RTS at 1733 GMT.

4/11/02

CAM/MR

Lowered the SE booms at 2003 GMT for 25M MFR maintenance.  There is no failure for this instrument.  Raised the booms and RTS at 2106 GMT.

4/2/02

CAM

The 60M tower SE booms needed to be lowered momentarily to reseat the MFR connections.  Lowered the SE booms at 1754 GMT.  Raised the booms and RTS at 1755 GMT.

11/01/01

MR

Vic Morris requested that the upwelling wiring be checked due to the readings on the US and UIR are bad. Lowered the tower at 2050 GMT.  T/S found that the 100K precision resistor between the CR10 and multiplexer was not installed.  Installed the resistor and notified the mentor of the change.  Raised the tower at 2120 GMT.

8/7/01-8/8/01

MR

Southeast booms lowered at 1918 GMT on 8/7 for 25M and 60M T/H calibrations.  Booms were left down at ground level overnight.  Raised booms and RTS at 1840 GMT on 8/8.

7/20/01

MR

The tower will be lowered due to work being performed on the PGS system.  No work will be performed on this instrument. Lowered tower at 1450 GMT.  Performed PGS system maintenance.  Raised tower at 1700 GMT.

7/13/01

MR

The 25m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2000 GMT.  Marc Fisher performed maintenance on the PGS system.  Raised the tower at 2130.  RTS at 2145 GMT.

7/12/01

MR

The 60m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2100 GMT.  Marc Fisher performed maintenance on the PGS system.  Raised the tower at 2135.  RTS at 2145 GMT.

1/18/00

KT

At 1905 GMT the GPS antenna located on the West end of the Optical Trailer was found hanging. At 1907 GMT reseated antenna. RTS at 1912 GMT.

1/6/2000

CW/CM

Site Ops lowered the 60M tower SE booms from 1941 GMT to 2123 GMT and installed the new GPS Array whip antenna at the 60M level. This antenna is the subject of BCR SGP-00256. Instrument mentor reported that by looking at the statistics of data collection before and after the antenna was replaced, it appears as if there is substantial improvement in the amount of data collected. Before the change, the system as a whole was collecting about 69% of the data. The day following installation, the collection percentage was 81%. Also, the stations where there was the most data loss had the highest increase in data collected. Here is a summary by site:

STATION BEFORE        AFTER CHANGE

-------                 ------                  -----                   ------

2N3W               81%                  87%                  +6%

2N2W               93                     95                     +2

2N0W               52                     80                     +28

2N1E                53                     73                     +20

1N3W               64                     75                     +11

1N2W               65                     78                     +13

1N1W               66                     82                     +16

1N0W               64                     76                     +12

0N2W               61                     77                     +16

0N1W               93                     96                     +3

0N0W               78                     80                     +2

1S3W               56                     73                     +17

1S1W               94                     95                     +1

1S0W               50                     70                     +20

 

TOTAL  69                     81                     +12

9/1/99

CW

Changed out the repeater antenna on the 60M tower with a yagi style antenna

 8/11/99

CW

GPS instrument array installed. The array consists of 14 stations that are located at various locations surrounding the Central Facility. The data from these instruments are transmitted back to a repeater installed on the 60M tower and then to the receiver mounted on the roof of the Optical trailer. The system PC is located inside the Optical trailer.